OpenStudy (anonymous):

f/g(x) = (x^2 +6) ---------- (2x^2-5) The Domain = ALL REAL NUMBERS - {2values} what are these 2 values? ____ (smaller value) ____ (larger value)

OpenStudy (anonymous):

So we are dividing two expressions, but when dividing, you want to make sure you dont divide by 0, because thats a math sin.

OpenStudy (anonymous):

i orginally thought the answers were 0 and 5

OpenStudy (anonymous):

So lets set the bottom polynomial equal to zero and solve for x: \[2x^{2} - 5 = 0 \Leftrightarrow 2x^{2} = 5\Leftrightarrow x^{2} = \frac{5}{2}\Leftrightarrow x = \pm \sqrt{\frac{5}{2}}\]

OpenStudy (anonymous):

0 and 5 are ok to plug into the funtion, because: \[2(0)^{2}-5 = 0-5 = -5\] and \[2(5)^{2}-5 = 50-5 = 45\] We wouldnt be dividing by 0 in those cases, so the division is ok, thats why 0 and 5 arent the answer

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

Basically, the rule is this: If you want to see what numbers to exclude from your domain, set the denominator equal to 0 and solve :)

OpenStudy (anonymous):

thank you! that makes it easier